OpenStudy (kirbykirby):

Verifying the property of the Gamma function: Gamma(n) = (n-1)!

OpenStudy (kirbykirby):

\[\Gamma (\alpha)=\int\limits_{0}^{\infty}x^{\alpha-1}e^{-x}dx\] I understand how to get to the relationship \[\Gamma (\alpha)=(\alpha-1)\Gamma (\alpha-1)\] for any alpha >1 But I'm not sure how t get to :\[\Gamma (n)=(n-1)!\] for n is a positive integer

OpenStudy (anonymous):

really? you have done all the hard work!

OpenStudy (kirbykirby):

I dunno why I can't see this.. my brain is fried probably from doing the first part LOL

OpenStudy (anonymous):

yeah i guess so

OpenStudy (anonymous):

\(n!=n(n-1)!\)

OpenStudy (anonymous):

that is \(n!\) can be defined as the recursion \(1!=1\) and \(n!=n(n-1)!\)

OpenStudy (kirbykirby):

OH i see right!! Oh thank you. Haha oh my that was fairly simple :P

OpenStudy (kirbykirby):

Thanks :)!!

OpenStudy (anonymous):

yeah way way more simple than showing \(\Gamma(\alpha)=(\alpha-1)\Gamma(\alpha -1)\)

OpenStudy (kirbykirby):

For sure!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

like opening the jar after someone already loosened it

OpenStudy (kirbykirby):

@satellite73 I couldn't help but wonder how do you keep your equations on the same line?